# entity resolution AL active learning

### v2

@@ Abstract.
er is a fundamental task in data integration that enables insights from heterogeneous data sources. The primary challenge of ER lies in classifying record pairs as matches or non-matches, which in multi-source entity resolution (ER) (MS-ER) scenarios can become complicated due to data source heterogeneity and scalability issues. Existing methods for MS-ER generally require labeled record pairs, and such methods fail to effectively reuse models across multiple ER tasks. We propose MoRER (Model Repositories for Entity Resolution), a novel method for building a model repository consisting of classification models that solve ER problems. By leveraging feature distribution analysis, MoRER clusters similar ER tasks, thereby enabling the effective initialization of a model repository with a moderate labeling effort. Experimental results on three multi-source datasets demonstrate that MoRER achieves comparable or better results to methods that have label-limited budgets, such as active learning and transfer learning approaches, while outperforming self-supervised approaches that utilize large pre-trained language models. When compared to supervised transformer-based methods, MoRER achieves comparable or better results, depending on the training data size. Importantly, MoRER is the first method for building a model repository for ER problems, facilitating the continuous integration of new data sources by reducing the need for generating new training data.
@@ 1. Introduction
Figure 1. Motivation of reusing solved ER tasks for new tasks. The data sources $D_{1}$ and $D_{2}$ are already linked utilizing similarity feature vectors and a model $M_{1,2}$ to label each record pair. The question is whether the derived model $M_{1,2}$ can also be applied to the new data source, $D_{3}$, to match it to $D_{1}$ and $D_{2}$, or if new models have to be generated.
@@
The quality of the ER result between two data sources is highly dependent on the classification step (Stonebraker and Ilyas, 2018). This step classifies record pairs into matches and non-matches based on attribute value similarities or derived embeddings from pre-trained language models. State-of-the-art methods use machine learning (ML) based approaches since they have been shown to achieve high-quality results (Doan et al., 2020). However, the majority of ML methods require labeled training data to build a classification model, where generating such data can be a time-consuming and expensive task (Mudgal et al., 2018a).
@@
Typical use cases for model repositories are found in research institutions and companies that require the continuous integration of new and diverse data sources.
@@
For example, in healthcare, hospitals and clinics often need to integrate patient records from various systems to create unified medical histories in order to enable better diagnosis, treatment, and research (Padmanabhan et al., 2019). Similarly, in census data analysis, government agencies must reconcile records from multiple surveys and administrative sources to ensure accurate population statistics and policy planning (Linacre et al., 2022). In e-commerce, product comparison portals and marketplaces integrate product catalogs from different vendors to provide search and recommendation functionalities for users.
@@
In all these scenarios, our method will allow organizations to reuse previously trained models from solved ER tasks, significantly reducing the labeling effort and computational costs. Furthermore, our method lays the groundwork for establishing ERmatching services, enabling users to solve any ER problem by leveraging existing models. This capability is particularly valuable in dynamic environments where new data sources are continuously added, ensuring scalability and efficiency in real-world applications.
@@
Fig. 1 illustrates the motivation of our proposed method. Assuming we have two integrated data sources, $D_{1}$ and $D_{2}$, consisting of products specified by the shown attributes. The integration of these data sources requires the solution of the ER problem given by the classification task of the similarity feature vectors into matches and non-matches. These similarity features are derived from the attribute-wise comparisons between record pairs. Therefore, the classification model, $M_{1,2}$, based on labeled record pairs, is trained and used to classify the record pairs $A1-B1,A2-B2,A3-B3,A3-B4\text{ and }A4-B1$. Later on, a new data source $D_{3}$ needs to be integrated, where we calculate similarity features between data sources $D_{1}$ and $D_{3}$ as well as between $D_{2}$ and $D_{3}$.
@@
In a naive solution, we would then build one new classification model, $M_{1,3}$ and $M_{2,3}$, for each new ER task between the data sources $D_{1}$ and $D_{3}$ and $D_{2}$ and $D_{3}$, respectively. Due to the increasing number of models with the growing number of integrated data sources, the labeling effort for generating training data would, however, increase substantially. Instead of building new models $M_{1,3}$ and $M_{2,3}$, we aim to reuse the existing model $M_{1,2}$ that has already solved the (potentially similar) ER task between $D_{1}$ and $D_{2}$. As the amount of data to be integrated increases, it becomes crucial to efficiently generate training data for similar ER problems to ensure effective and scalable classification. Our proposed method, MoRER, addresses these challenges.
@@
An alternative approach would be to unify all existing ER problems to create a comprehensive training dataset. However, individual ER problems can have different distributions of their similarity features. This would make it challenging for a single classification model to account for the potentially diverse similarity distributions. To overcome such an issue, the idea would be to create new training data for each new ER problem and combine it with the existing one. However, such an approach is not scalable, since it requires the training data generation for each new ER problem. Fig. 2 illustrates the variations in similarity distributions calculated by the Jaccard function on the title for products on the WDC computer dataset (Primpeli et al., 2019), which consists of four data sources. Given these very diverse similarity distributions, a single unified classification model would struggle to capture the diverse characteristics of matches and non-matches. In addition to the different feature spaces, new ER problems might not fit into an existing model, so retraining is required.
@@
Therefore, we suggest a method to create an ER model repository and to reuse the maintained models for similar ER problems. In this paper, we make the following contributions:
@@
  • We propose a novel method to build an ER model repository of classification models for ER problems. Our method analyzes the similarity feature space to determine the most appropriate model for an unsolved novel ER problem.
@@
  • Our proposed method efficiently initializes the repository with models that require only a little labeling effort. To do so, we build an ER problem graph and cluster ER tasks according to their similar feature spaces. For each cluster, we build a classification model using the most relevant similarity feature vectors as training data.
@@
The remainder of this paper is organized as follows. In Section 2 we formalize our problem, and in Section 3 we review related work on label-efficient and state-of-the-art ER methods. In Section 4 we then describe our method in detail, which we evaluate in Section 5 and compare it with several state-of-the-art approaches. In Section 6, we discuss our results and provide recommendations. We conclude in Section 7 with future directions.
@@
Figure 2. Example of the similarity distributions using the Jaccard similarity for title considering the ER problems in the WDC-computer data set. Each of the five lines of different colors represents an ER problem.
@@ 2. Problem Definition
Following the notation given in Table 1, given a set $\mathcal{D}$ of data sources $\{D_{1},\ldots,D_{z}\}$. An ER problem $p_{k,l}$ is defined for a pair of data sources $D_{k}$ and $D_{l}$ consisting of similarity features for each record pair. We utilize an initial set $\mathcal{P_{I}}$ of ER problems $p_{i,l}$ to set up the ER model repository. We aim to solve new ER problems $p_{i,l}\in\mathcal{P_{U}}$ utilizing the initial ER problems $\mathcal{P_{I}}$.
An ER problem $p_{i,l}$ is defined for a data source pair $D_{i},D_{l}\in\mathcal{D}$ and consists of a set of similarity feature vectors $w$. Each similarity feature vector $w\in p_{k,l}$ represents a record pair, where each entry in $w$ corresponds to a similarity in the range of [0,1], calculated based on attribute value comparisons using a similarity function. We denote an entry of a similarity feature vector $w$ as feature $f$. We assume that even for heterogeneous data sources, common features $f$ exist based on common attributes between the different data sources, which are significant to describe the entities of a certain domain, such as the title of products, price information, etc. In the example in Fig. 1, $\mathcal{P_{I}}$ consists of the ER problem $P_{D_{1},D_{2}}$ and $\mathcal{P_{U}}$ includes the ER problems $P_{D_{2},D_{3}}$ and $P_{D_{1},D_{3}}$.
To reduce the effort of building new classification models and the underlying label generation process, our solution groups similar ER problems $p_{k,l}\in\mathcal{P_{I}}$. We determine the similarity between pairs of ER problems utilizing univariate distribution tests to group similar problems into clusters $\mathcal{C_{P}}$. We hypothesize that the ER tasks within a single cluster $C\in\mathcal{C_{P}}$ can be effectively classified using a unified classification model $M_{C}$.
The analysis utilizes the similarity distribution $d^{f}_{k,l}$ and the cumulative distribution function $CDF^{f}_{k,l}$ for each feature $f$ based on a data source pair $D_{k}$ and $D_{l}$. We construct an ER problem similarity graph $G_{P}=(\mathcal{P_{I}\cup\mathcal{P_{U}}},E)$ where the edges represent the determined similarities between ER problems. The graph is used for clustering and is extendable to integrate new ER problems of $\mathcal{P_{U}}$.
@@
We define the problem of solving ER problems $p_{D_{x},D_{z}}\in\mathcal{P_{U}}$ as the search for a similar cluster $C\in\mathcal{C_{P}}$ to apply the corresponding model $M_{C}$ on the ER problem between $D_{x}$ and $D_{z}$. Moreover, we update the determined model $M_{C}$ if the ER problems of cluster $C$ become dissimilar or the coverage of training data of a certain cluster decreases below a certain threshold.
@@ 3. Related Work
In this section, we discuss state-of-the-art ER approaches from traditional ML-based methods to language model-based solutions such as utilizing LLMs. In general, the main goal of our method is to reduce the effort for the model generation step. Therefore, we discuss budget-limited machine learning (ML) methods in ER, such as self-supervised, active, and transfer learning.
Entity Resolution. ER methods classify record pairs based on features derived from similarity functions (Christen, 2012). In the past, supervised ER techniques have been developed, leveraging both traditional ML algorithms (Konda et al., 2016; Köpcke et al., 2010; Bilenko and Mooney, 2003) and deep learning-based methods (Mudgal et al., 2018b; Ebraheem et al., 2017). The advent of transformer-based language models (Vaswani et al., 2017) has further advanced ER by enabling the representation of textual and unstructured data as embeddings, which can be directly compared (Li et al., 2020; Brunner and Stockinger, 2020). Transformers, powered by self-attention mechanisms, process input data in parallel, making them highly effective for tasks like natural language processing and sequence modeling. State-of-the-art methods, such as Ditto  (Li et al., 2020), fine-tune transformer models like BERT (Devlin et al., 2019) and DistilBERT (Sanh et al., 2019) to classify record pairs as matches or non-matches. Auto-EM (Zhao and He, 2019) relies on pre-trained models for different entity types to classify value pairs as match or non-match. For a new linkage problem, the relevant entity types are identified, allowing the corresponding models to be utilized for classifying the record pairs. MultiEM  (Zeng et al., 2024) clusters records from multiple data sources utilizing embeddings from pre-trained language models such as BERT. The approach follows a hierarchical merge strategy where two data sources are compared to avoid the comparison of all data source pairs. While the proposed approach shows its strength in efficiency, it requires the identification of effective hyperparameters, the similarity threshold $m$, and clustering-related parameters.
Peeters et al. (Peeters and Bizer, 2023) showed the feasibility of using large language models for ER on various real-world datasets and different prompt variations without any fine-tuning. Recent work (Zhang et al., 2025) achieved comparable results by using problem-dependent fine-tuned small language models (Zhang et al., 2025) with a significant cost reduction.
@@
Active Learning. AL approaches aim to minimize the labeling effort that is particularly important for multi-source ER due to the numerous classification tasks involved. Most AL methods focus on pairwise ER problems (Mozafari et al., 2014; Christen et al., 2019; Wang et al., 2015; Ngonga Ngomo and Lyko, 2012) utilizing uncertainty measures and distances between feature vectors. Only one study (Primpeli and Bizer, 2021) has specifically addressed multi-source ER by leveraging similarity graphs generated from the ER process. This work employs the concept of transitive closure to identify informative links, enabling the detection of potential false negatives and false positives. Specifically, it assesses whether two records, classified as non-matches, are connected through the transitive closure. Potential false positives are identified using the minimum cut of the graph. At the same time, the resulting graph also facilitates the identification of false negatives, represented as missing edges among record pairs within connected components.
@@
Limitations. Despite substantial progress in ER through machine learning, active learning, and transfer learning, existing approaches face several significant limitations, particularly in multi-source and dynamic integration settings. Most existing approaches focus on pairwise ER, requiring the generation of new classification models for each ER problem when the similarity feature space changes, resulting in a high training effort. In the following, we summarize the specific limitations for each area.
  • Language Model. While pretrained language model-based and LLM methods have substantially advanced ER performance, they struggle with practical constraints. Fine-tuning such models requires large labeled datasets, making them costly and impractical in scenarios with limited annotation resources. Moreover, LLM approaches are monetarily intensive and struggle to scale to the massive number of record pairs present in real-world multi-source ER tasks.
  • Active Learning. These techniques have been proposed to reduce labeling effort by selecting the most informative instances for annotation. However, most AL approaches are tailored to pairwise ER and do not generalize to the complexities of multi-source settings, where the number of possible classification tasks grows rapidly. Even recent multi-source AL methods, such as Almser, focus on optimizing label selection within known tasks but lack mechanisms to determine when existing models can be reused or when new models (and thus fresh training data) are necessary for new ER problems.
  • Transfer learning seeks to reuse models or knowledge from previously solved ER tasks. However, in practice, these methods are limited by their inability to effectively select the most appropriate source task(s) when multiple candidates exist. Most transfer learning approaches require significant similarity between source and target domains, which is difficult to guarantee in heterogeneous and evolving multi-source environments. Furthermore, existing methods do not provide systematic procedures for evaluating model suitability.

### v3

@@ Abstract.
er is a fundamental task in data integration that enables insights from heterogeneous data sources. The primary challenge of ER lies in classifying record pairs as matches or non-matches, which in multi-source entity resolution (ER) (MS-ER) scenarios can become complicated due to data source heterogeneity and scalability issues. Existing methods for MS-ER generally require labeled record pairs, and such methods fail to effectively reuse models across multiple ER tasks. We propose MoRER (Model Repositories for Entity Resolution), a novel method for building a model repository consisting of classification models that solve ER problems. By leveraging feature distribution analysis, MoRER clusters similar ER tasks, thereby enabling the effective initialization of a model repository with a moderate labeling effort. Experimental results on three multi-source datasets demonstrate that MoRER achieves comparable or better results to methods that have label-limited budgets, such as active learning and transfer learning approaches, while outperforming self-supervised approaches that utilize large pre-trained language models. When compared to supervised transformer-based methods, MoRER achieves comparable or better results, depending on the size of the training data set used.
Record Linkage, Model Reuse, Distribution Analysis
@@ 1. Introduction
Figure 1. Motivation of reusing solved ER tasks for new tasks. The data sources $D_{1}$ and $D_{2}$ are already linked utilizing similarity feature vectors and a model $M_{1,2}$ to label each record pair. The question is whether the derived model $M_{1,2}$ can also be applied to the new data source, $D_{3}$, to match $D_{3}$ to $D_{1}$ and $D_{2}$, or if new models have to be generated.
@@
The quality of the ER result between two data sources is highly dependent on the classification step (Stonebraker and Ilyas, 2018). This step classifies record pairs into matches and non-matches based on attribute value similarities or derived embeddings from pre-trained language models. State-of-the-art methods use machine learning (ML) based approaches since they have been shown to achieve high-quality results (Doan et al., 2020). However, the majority of ML methods require labeled training data to build a classification model, where generating such data can be a time-consuming and expensive task (Mudgal et al., 2018).
@@
Typical use cases for model repositories are found in research institutions and companies that require the continuous integration of new and diverse data sources. For example, in healthcare, hospitals and clinics often need to integrate patient records from various systems to create unified medical histories in order to enable better diagnosis, treatment, and research (Padmanabhan et al., 2019). Similarly, in census data analysis, government agencies must reconcile records from multiple surveys and administrative sources to ensure accurate population statistics and policy planning (Linacre et al., 2022). In e-commerce, product comparison portals and marketplaces integrate product catalogs from different vendors to provide search and recommendation functionalities for users.
@@
In all these scenarios, our method will allow organizations to reuse previously trained models from solved ER tasks, significantly reducing both labeling efforts and computational costs. Furthermore, our method lays the groundwork for establishing ER matching services, enabling users to solve any ER problem by leveraging existing models. This capability is particularly valuable in dynamic environments where new data sources are continuously added, ensuring scalability and efficiency in real-world applications.
@@
Fig. 1 illustrates the motivation of our proposed method. Assuming we have two integrated data sources, $D_{1}$ and $D_{2}$, consisting of products specified by the shown attributes. The integration of these data sources requires the solution of the ER problem given by the classification task of the similarity feature vectors into matches and non-matches. These similarity features are derived from the attribute-wise comparisons between record pairs. Therefore, the classification model, $M_{1,2}$, based on labeled record pairs, is trained and used to classify the record pairs $A1-B1,A2-B2,A3-B3,A3-B4\text{ and }A4-B1$. Later on, a new data source $D_{3}$ needs to be integrated, where we calculate similarity features between data sources $D_{1}$ and $D_{3}$, as well as between $D_{2}$ and $D_{3}$.
@@
In a naive solution, we would build one new classification model, $M_{1,3}$ and $M_{2,3}$, for each new ER task between the data sources $D_{1}$ and $D_{3}$ and $D_{2}$ and $D_{3}$, respectively. Due to the increasing number of models with the growing number of integrated data sources, the labeling effort for generating training data would, however, increase substantially. Instead of building the new models $M_{1,3}$ and $M_{2,3}$, our aim is to reuse the existing model $M_{1,2}$ that has already solved the (potentially similar) ER task between $D_{1}$ and $D_{2}$. As the amount of data to be integrated increases, it becomes crucial to efficiently generate training data for similar ER problems to ensure effective and scalable classification. Our proposed method, MoRER, addresses these challenges.
@@
An alternative approach would be to unify all existing ER problems to create a comprehensive training dataset. However, individual ER problems can have different distributions of their similarity features. This would make it challenging for a single classification model to account for the potentially diverse similarity distributions. To overcome this issue, the idea would be to create new training data for each new ER problem and combine these with existing one. However, such an approach is not scalable, since it requires the training data generation for each new ER problem. Fig. 2 illustrates the variations in similarity distributions calculated by the Jaccard string comparison function (Christen, 2012) on the title for products on the WDC computer dataset (Primpeli et al., 2019), which consists of four data sources. Given these very diverse similarity distributions, a single unified classification model would struggle to capture the diverse characteristics of matches and non-matches. Moreover, new ER problems might not fit into an existing model, so retraining is required.
To the best of our knowledge, no current approach takes model reuse into account. Existing approaches assume that training data are available to train a model and that this model can be used for all future record pairs, as in progressive or incremental ER (Gazzarri and Herschel, 2023).
@@
Therefore, we suggest a method to create an ER model repository and to reuse the maintained models for similar future ER problems. The reuse of models is highly relevant where new data sources must be integrated in target data sources such as data warehouses or for ad hoc analysis with data from data lakehouses. In this paper, we make the following contributions:
@@
  • We propose a novel method to build an ER model repository of classification models for ER problems. Our method analyzes the similarity feature space to determine the most appropriate model for an unsolved novel ER problem. Unlike existing ER methods, our method distinguishes between ER tasks with different similarity feature spaces and selects the appropriate models for the various ER tasks.
@@
  • Our proposed method efficiently initializes the repository with models that require only little labeling effort. To do so, we build an ER problem graph and cluster ER tasks according to their similar feature spaces. For each cluster, we build a classification model using the most relevant similarity feature vectors as training data. The resulting clusters and models can be stored in a backend system and used for future ER tasks.
@@
The remainder of this paper is organized as follows. In Section 2 we formalize the problem we aim to solve, and in Section 3 we review related work on label-efficient and state-of-the-art ER methods. In Section 4 we then describe our method in detail, which we evaluate in Section 5 and compare it with several state-of-the-art approaches. In Section 6, we discuss our results and provide recommendations. We conclude in Section 7 with future research directions.
@@
Figure 2. Example of the similarity distributions calculated using the Jaccard string comparison function (Christen, 2012) for title considering the ER problems in the WDC computer data set. Each of the five lines represents an ER problem.
@@ 2. Problem Definition
Following the notation shown in Table 1, we are given a set $\mathcal{D}$ of data sources $\{D_{1},\ldots,D_{z}\}$. An ER problem $p_{k,l}$ is defined for a pair of data sources $D_{k}$ and $D_{l}$ consisting of similarity features for each record pair. We utilize an initial set $\mathcal{P_{I}}$ of ER problems $p_{i,l}$ to set up the ER model repository. We aim to solve new ER problems $p_{i,l}\in\mathcal{P_{U}}$ utilizing the initial ER problems $\mathcal{P_{I}}$.
An ER problem $p_{i,l}$ is defined for a data source pair $D_{i},D_{l}\in\mathcal{D}$ and consists of a set of similarity feature vectors $w$. Each similarity feature vector $w\in p_{k,l}$ represents a record pair, where each element of $w$ corresponds to a similarity in the range of [0,1], calculated based on attribute value comparisons using a similarity function. We denote an element of a similarity feature vector $w$ as feature $f$. We assume that even for heterogeneous data sources, common features $f$ exist based on common attributes between the different data sources, which are significant to describe the entities of a certain domain, such as the title of products, price information, etc. In the example in Fig. 1, $\mathcal{P_{I}}$ consists of the ER problem $P_{D_{1},D_{2}}$ and $\mathcal{P_{U}}$ includes the ER problems $P_{D_{2},D_{3}}$ and $P_{D_{1},D_{3}}$.
To reduce the effort of building new classification models and the underlying label generation process, our solution groups similar ER problems $p_{k,l}\in\mathcal{P_{I}}$. We determine the similarity between pairs of ER problems utilizing univariate distribution tests to group similar problems into clusters $\mathcal{C_{P}}$. We hypothesize that the ER tasks within a certain cluster $C\in\mathcal{C_{P}}$ can be effectively classified using a unified classification model $M_{C}$.
The analysis uses the similarity distribution $d^{f}_{k,l}$ and the cumulative distribution function $CDF^{f}_{k,l}$ for each feature $f$ based on a data source pair $D_{k}$ and $D_{l}$. We construct an ER problem similarity graph $G_{P}=(\mathcal{P_{I}\cup\mathcal{P_{U}}},E)$ where the edges represent the determined similarities between ER problems. The graph is used for clustering and is extendable to integrate new ER problems of $\mathcal{P_{U}}$.
@@
We define the problem of solving ER problems $p_{D_{x},D_{z}}\in\mathcal{P_{U}}$ as the search for a similar cluster $C\in\mathcal{C_{P}}$ to apply the model $M_{C}$ on the ER problem between $D_{x}$ and $D_{z}$. Moreover, we update the identified model $M_{C}$ if the ER problems of cluster $C$ become dissimilar or the coverage of training data of a certain cluster decreases below a certain threshold.
@@ 3. Related Work
In this section, we discuss state-of-the-art ER approaches from traditional machine learning (ML) based methods to language model based solutions such as utilizing LLMs. In general, the main goal of our method is to reduce the effort for the model generation step. Therefore, we discuss budget-limited ML methods in ER, such as self-supervised, active, and transfer learning.
[…]
er methods classify record pairs based on features derived from similarity functions (Christen, 2012). In the past, supervised ER techniques have been developed, leveraging both traditional ML algorithms (Konda et al., 2016; Köpcke et al., 2010; Bilenko and Mooney, 2003) and deep learning-based methods (Mudgal et al., 2018; Ebraheem et al., 2017). The advent of transformer-based language models (Vaswani et al., 2017) has further advanced ER by enabling the representation of textual and unstructured data as embeddings, which can be directly compared (Li et al., 2020; Brunner and Stockinger, 2020). Transformers, powered by self-attention mechanisms, process input data in parallel, making them highly effective for tasks like natural language processing and sequence modeling. State-of-the-art methods, such as Ditto  (Li et al., 2020), fine-tune transformer models like BERT (Devlin et al., 2019) and DistilBERT (Sanh et al., 2019) to classify record pairs as matches or non-matches. Auto-EM (Zhao and He, 2019) relies on pre-trained models for different entity types to classify value pairs as match or non-match. For a new ER problem, the relevant entity types are identified, allowing the corresponding models to be utilized for classifying the record pairs. Unicorn  (Fan et al., 2024) generates an universal model addressing various matching tasks such as entity alignment, ontology matching, or ER, by using a unified encoder model and adopting a mixture-of-experts model. MultiEM  (Zeng et al., 2024) clusters records from multiple data sources utilizing embeddings from pre-trained language models. This approach follows a hierarchical merge strategy where two data sources are compared to avoid the comparison of all data source pairs. While the proposed approach shows its strength in efficiency, it requires the identification of effective hyperparameters, the similarity threshold $m$, and clustering-related parameters. Peeters et al. (Peeters and Bizer, 2023) showed the feasibility of using large language models for ER on various real-world datasets and different prompt variations without any fine-tuning. Recent work (Zhang et al., 2025) achieved comparable results by using problem-dependent fine-tuned small language models (Zhang et al., 2025) with a significant cost reduction.
@@
Active Learning. AL approaches aim to minimize the labeling effort that is particularly important for multi-source ER due to the numerous classification tasks involved. Most AL methods focus on pairwise ER problems (Mozafari et al., 2014; Christen et al., 2019; Wang et al., 2015; Ngonga Ngomo and Lyko, 2012) utilizing uncertainty measures and distances between feature vectors. Only one study (Primpeli and Bizer, 2021) (Almser) has specifically addressed multi-source ER by leveraging similarity graphs generated from the ER process. This work employs the concept of transitive closure to identify informative links, enabling the detection of potential false negatives and false positives. Specifically, it assesses whether two records, classified as non-matches, are connected through the transitive closure. Potential false positives are identified using the minimum cut of the graph. At the same time, the resulting graph also facilitates the identification of false negatives, represented as missing edges among record pairs within connected components.
@@
Limitations Despite substantial progress in ER through machine learning, active learning, and transfer learning, existing approaches face several significant limitations, particularly in multi-source and dynamic integration settings. Most existing approaches focus on pairwise ER, requiring the generation of new classification models for each ER problem when the similarity feature space changes, resulting in a high training effort. In the following, we summarize the specific limitations for each area.
Most AL approaches are tailored to pairwise ER and do not generalize to the complexities of multi-source settings, where the number of possible classification tasks grows rapidly. Even recent multi-source AL methods, such as Almser (Primpeli and Bizer, 2021), focus on optimizing label selection within known tasks but lack mechanisms to determine when existing models can be reused or when new models (and thus fresh training data) are necessary for new ER problems.
Transfer learning methods are limited by their inability to effectively select the most appropriate source task(s) when multiple candidates exist. Most transfer learning approaches require significant similarity between source and target domains, which is difficult to guarantee in heterogeneous and evolving multi-source environments. Furthermore, existing methods do not provide systematic procedures for evaluating model suitability.
While pretrained language model-based and LLM methods have substantially advanced ER performance, they struggle with practical constraints. Fine-tuning such models requires large labeled datasets, making them costly and impractical in scenarios with limited annotation resources. Moreover, LLM approaches are monetarily intensive and struggle to scale to the massive number of record pairs present in real-world multi-source ER tasks.